OpenStudy (anonymous):

Simplify the following:

OpenStudy (anonymous):

\[e ^{\frac{ 1 }{ 3 }}logA\] the 1/3*logA all indicies.

OpenStudy (anonymous):

Also logA+logB-2logC as a single fraction.

OpenStudy (unklerhaukus):

1.\[e^{\frac13\log A}\]

OpenStudy (anonymous):

is log here the same as ln?

OpenStudy (unklerhaukus):

\(\log x\) with out a base written is usually means logarithm to base ten ; \(\log x=\log_{10} x\) \(\ln x\) is logarithm to the natural base \(e\) ; \(\ln x=\log_e x\)

geerky42 (geerky42):

The funny thing is that Wolfram Alpha uses log x as natural logarithm of x. log can represent as natural log.

OpenStudy (unklerhaukus):

Apply some logarithm laws \[\boxed{n\log_b x=\log_bx^n}\]\[\boxed{b^{\log_b x}=x}\] \[\boxed{\log_b x=\frac{\log_c x}{ \log_c b}}\] \[\boxed{\log_b x+\log_by=\log_bxy}\]\[\boxed{\log_b x-\log_by=\log_b\frac xy}\]

OpenStudy (unklerhaukus):

i dont know why wolfram breaks the convention with logarithms but it sure can be confusing @geerky42